HUBBLE SPACE TELESCOPE DAILY REPORT #846 PERIOD COVERED: 0800 EST Daily Status Report as of 077/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WFC 4658 (Determination of Extragalactic Distance Scale) The Wide Field Camera was used to make two images of sections of the galaxy M-101 (aka, NGC-5457) to look for Cepheid variable stars. The baseline guide star acquisition failed due to an attitude error. An Attitude Reference Update/ Planned Real-Time Slew (ARU/PRT) was uplinked to the spacecraft. The exposures were made in course track. During the first exposure the gyros were commanded into High Mode in real-time. A loss of lock occurred about five times. The gyros were then commanded back to Low Mode. When the image was displayed it was smeared. The second image was reported to look okay. 1.2 Completed FOC/96/WFC 3325 (Structure of Globular Clusters) The Faint Object Camera and the Wide Field Camera (f/96) were both used in parallel to make images of the giant globular cluster Omega Centauri (aka, NGC-5139). All of the FOC and all of the WFPC images were "excellent". 1.3 Initiated GHRS/2 3755 (Absorption Cross-section of Nearby Galaxies) The High Resolution Spectrograph was used to initiate observations of quasar 1543-489 which will be used to probe for absorption lines produced by the halos of foreground galaxies. The first acquisition was successful, and the target had a flux of 544. During the second observation there was six losses of lock while the gyros were being run in the High Rate Mode. Lock was recovered each time. The observations are still ongoing, and will be analyzed later. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 2 Successful : 1 (Failure at 076/1225Z, See Note 1.1) Scheduled Reacquisitions : 12 Successful : 12 Losses of Lock : At least 2 times (See Note 1.1, and 1.3) 2.2 FHST Updates: Scheduled : 6 Successful : 5 (Failure at 076/2151Z) 2.3 Significant HSTARs: None 2.4 Operations Note: None 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete GHRS/2 3755 (Absorption Cross-Section of Nearby Galaxies) 3.2 Complete Three Sets of FOC/96 3325 (UV Imaging of Nearby Galaxies With FOC/96) 3.3 Complete GHRS/2 1177 (Search For Hot Plasmas in Outer Atmosphere) 3.4 Uplink Updated Ephemeris Tables, Per MOSES 2871 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: FGS #2 Anomaly Investigation Further testing and analysis of the FGS #2 anomalous behavior are still ongoing. In the interim FGS #2 will not be used for observations. 5.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. -END-